OpenStudy (anonymous):

Josh has 12 coins, he has four times as many dimes as quarters the total value is 1.00. How many dimes nickles and quarters does he have?

OpenStudy (imstuck):

You have to solve this using a system of equations, one that deals with the NUMBER of coins and another that deals with the value of the coins. He has 12 coins total, 4 times as many dimes as quarters. That will be the first equation we will have to write.

OpenStudy (anonymous):

4x=d

OpenStudy (anonymous):

and x=q

OpenStudy (anonymous):

Not sure where to go from there

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

Sorry, I had to tend to something...I'm here. give me a sec, ok?

OpenStudy (imstuck):

I'm here, just give me a sec...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@IMStuck You back?

OpenStudy (larseighner):

There are 12 coins, nickels, dimes and quarters: n + d + q = 12 There are 4 times as many dimes as quarters: 4q = d The total values is 100 (just for simplicity's sake, using cents instead of dollars): 5n + 10d +25q = 100 There are three equations in three unknowns. Solve simultaneously.

OpenStudy (imstuck):

OMG!!!!! I totally overlooked the fact that there were nickels in this problem, too! No Wonder it wasn't working out! Good God AlmightY!!!!!!!

OpenStudy (imstuck):

Did you figure this out, Toasterino? Can you solve those 3 equations simultaneously?

OpenStudy (anonymous):

No, the 5n+10d+25q is what is confusing me, if i solve each they will all get 100

OpenStudy (imstuck):

Here are your equations, like LarsEighner said, ok: n+d+q=12 (that is the number of nickels + number of dimes + number of quarters) 4q=d (because there are 4 times as many dimes as quarters) 5n+10d+25q=100 (because a nickel is 5 cents, a dime is 10 cents and a quarter is 25) Start with the fact that d=4q and fill "4q" in for "d" in our "money equation" like this:

OpenStudy (imstuck):

5n+10(4d)+25q=100...5n+40q+25q=100...5n+65q=100. That is our first equation in 2 terms. Now we will do the same with the equation that relates the number of coins to each other, keeping in mind the fact that d=4q:

OpenStudy (imstuck):

n+d+q=12...n+4q+q=12...n+5q=12. That's our second equation in terms of two variables. So let's look at those 2 together, now:

OpenStudy (imstuck):

From above we have 5n + 65q = 100 and then n+ 5q = 12. We need to solve those for either n or q. I picked q, so let's do it like this:

OpenStudy (imstuck):

5n+65q=100 -5(n+5q=12)... 5n+65q=100 -5n -25q=-60 The n's cancel each other out (5n - 5n = 0n, right?) so we are left with 40q=40 (65q-25q=40q, and 100-60=40) 40q=40...q=1. So there is 1 quarter.

OpenStudy (imstuck):

Going back to one of our equations in two terms, we can sub in 1 for q and solve for n like this:

OpenStudy (imstuck):

4q=d...4(1)=d...d=4. So there are 4 dimes.

OpenStudy (imstuck):

NOW, going back to our original problem that says n+d+q=12, we can fill in the fact that q=1, and d=4 to solve for n. Like this:

OpenStudy (imstuck):

n+q+d=12...n+1+4=12...n+5=12...12-5=n...n=7.

OpenStudy (imstuck):

When you fill all that back into our money equation, we have 7(.05)+4(.10)+1(.25)=1.00... (there are 7 nickels and nickels are 5 cents each; there are 4 dimes and dimes are 10 cents each; there is one quarter and they are 25 cents each)... .35+.40+.25=1.00 See that?

OpenStudy (anonymous):

@IMStuck I understand it now, what was confusing was the equation 5n+10d+25q but now I understand that I had to substitute thanks for everything! As well as @LarsEighner for providing the fundamentals

OpenStudy (larseighner):

With a simple problem like this you can go it intuitively. There cannot be 3 quarters because then there would have to be 12 dimes, but that is more than 12 coins total. So there must be 1 or 2 quarters (of there were 0 quarters, you would have 12 nickels = 100, which cannot be). Now if there are 2 quarters, there must be 8 dimes, but that 50 + 80 which is > 100. So there must be 1 quarter. Then you know there are 4 dimes. That is 65 cents, leaving 35 cents in nickels which is 7 nickels. 7 nickels, 4 dimes, 1 quarter = 12 coins..

OpenStudy (imstuck):

Yay for all of us, then! ; )